# The prescribed metric on the boundary of convex subsets of anti-de Sitter space with a quasi-circle as ideal boundary

### v1

@@ 4.1. Definition of gluing maps
From now until the end of the paper, we will always assume that if $\Omega$ is a globally hyperbolic convex subset, then the induced metrics on its boundary components ±$\partial^{\pm}\Omega$ are isometric to $(\mathbb{D},h^{\pm})$, where $h^{\pm}$ have bounded derivatives of any order on the disc $\mathbb{D}$.
Note that from the proof of Lemma 4.4, the maps $\Pi_{l}^{\pm}\circ V^{\pm}$ extend to a homeomorphism $\partial(\Pi_{l}^{\pm}\circ V^{\pm}):\partial\mathbb{D}\to\partial_{\infty}\mathbb{H}^{2}$.
We define $\partial V^{\pm}:=\pi_{l}^{-1}\circ\partial(\Pi_{l}^{\pm}\circ V^{\pm})$, where $\pi_{l}$ is the extension of $\Pi_{l}$ to the boundary at infinity.
@@
Note that the isometries $V^{\pm}$ and the gluing map $\Phi_{\Omega}$ are defined up to composition by Möbius maps. Also note that Möbius maps do not necessarily preserve the metrics $h^{\pm}$. However, we still have a uniquely defined normalized gluing map.
@@
From the discussion above, in particular we get the following proposition
@@
4.2. Properness of the gluing maps
In this section we prove the properness of the gluing maps.
@@ 4.2. Properness of the gluing maps
Let $\epsilon>0$ and $(M_{p})_{p\in\mathbb{N}}$ a sequence of positive numbers. For any $k>1$ there exists $k^{\prime}>1$ such that for any globally hyperbolic convex subset $\Omega$ in which the induced metrics on ±$\partial^{\pm}\Omega$ have curvatures in $(-\frac{1}{\epsilon},-1-\epsilon)$ and any derivative of the metric of order $p$ is bounded by $M_{p}$. We assume that the gluing map is normalized. If the gluing map $\Phi_{\Omega}$ is $k$-quasi-symmetric then $\partial_{\infty}\Omega$ is a $k^{\prime}$ quasi-circle.
@@
[BDMS21, Lemma 8.4] Let $P$ be a totally geodesic spacelike plane. We denote by $P^{+}$ (resp $P^{-}$) the union of all future-oriented (resp past-oriented ) timelike geodesic segments of length $\frac{\pi}{2}$ starting orthogonally from $P$.
  • let $S\subset\mathbb{ADS}^{2,1}$ be a spacelike past convex surface, and let $P$ be a spacelike totally geodesic plane. In the neighborhood of all points $x\in S\cap P$ , the intersection $S\cap P^{+}$ is locally convex in the induced metric on $S$.
  • Let $S\subset\mathbb{ADS}^{2,1}$ be a spacelike future convex surface, and let $P$ be a spacelike totally geodesic plane. In the neighborhood of all points $x\in S\cap P$ , the intersection $S\cap P^{-}$ is locally convex in the induced metric on $S$.
@@
We argue by contradiction. Suppose that $\Omega_{n}$ is a sequence of normalized globally hyperbolic convex subsets such that $k^{\prime}_{n}$, the optimal quasi-symmetric constant of their ideal boundary $\partial_{\infty}\Omega_{n}$, converges to $\infty$. We will prove that $k_{n}$, the quasi-symmetric constants of their gluing maps $\Phi_{\Omega_{n}}$, must also go to $\infty$. Let $V^{\pm}_{n}:(\mathbb{D},h^{\pm}_{n})\to\partial^{\pm}\Omega_{n}$ be isometries such that the normalized gluing map $\Phi_{\Omega_{n}}=(\partial V^{-}_{n})^{-1}\circ(\partial V^{+}_{n})$. By applying an isometry of $\mathbb{ADS}^{2,1}$ to $\Omega_{n}$, we can assume that $\partial V^{+}_{n}(p)=(p,p)$ for any $p=0,1,\infty$. Since we assumed that the gluing map is normalized, it follows that $\partial V^{-}_{n}(p)=(p,p)$ for any $p=0,1,\infty$. Given $k^{\prime}_{n}\to\infty$, Proposition 2.5 implies that the width $w(C_{n})$ of $C_{n}:=\partial_{\infty}\Omega_{n}$ goes to $\frac{\pi}{2}$. After adjusting by isometries, we can assume that $C_{n}$ converges to a rhombus $C_{\diamond}$ as in [BDMS21, Example 6.7] (note that even after applying these isometries, we can still assume that $C_{n}$ and $C_{\diamond}$ are normalized).
@@
Let’s work in the projective model of $\mathbb{ADS}^{2,1}$ with the coordinates so that in the affine chart $x_{4}=1$, $\partial_{\infty}\mathbb{ADS}^{2,1}$ is the hyperboloid $x_{1}^{2}+x_{2}^{2}=x_{3}^{2}+1$, with $\mathbb{ADS}^{2,1}$ seen as the region $x_{1}^{2}+x_{2}^{2}<x_{3}^{2}+1$. We may then arrange that the vertices of $C_{\diamond}$ are the points (±$(\pm\sqrt{2},0,-1)$ and $(0,\pm\sqrt{2},1)$. Since $\partial^{+}\Omega_{n}$ is in the future of $\rm{CH}(C_{n})$ but contained in the invisible domain $E(C_{n})$, and since both $\overline{E(C_{n})}$ and $\rm{CH}(C_{n})$ converge to $\rm{CH}(C_{\diamond})$, we have that $\partial^{+}\Omega_{n}$ converge to $S^{+}_{\infty}$, where $S^{+}_{\infty}$ is the union of the two future faces of $\rm{CH}(C_{\diamond})$. Similarly, $\partial^{-}\Omega_{n}$ converge to $S^{-}_{\infty}$, where $S^{-}_{\infty}$ is the union of the two past faces of $\rm{CH}(C_{\diamond})$.
@@
Let $P$ be the spacelike totally geodesic plane equal to the intersection with $\mathbb{ADS}^{2,1}$ of the plane $z=0$ in $\mathbb{R}^{3}$. Let $a=\left(-\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},0\right)$, $b=\left(-\frac{\sqrt{2}}{2},-\frac{\sqrt{2}}{2},0\right)$, $c=\left(\frac{\sqrt{2}}{2},-\frac{\sqrt{2}}{2},0\right)$, and $d=\left(\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},0\right)$. Thus, $a,b,c,d$ are the intersection points of $\partial P$ with $C_{\diamond}$ occurring in the cyclic order $a,b,c,d$.
@@
For all $n\in\mathbb{N}$, let $a_{n},b_{n},c_{n},d_{n}$ be the intersection points of $C_{n}$ with $P$, such that $a_{n}\to a$, $b_{n}\to b$, $c_{n}\to c$, $d_{n}\to d$. Define $a_{n}^{\pm},b_{n}^{\pm},c_{n}^{\pm},d_{n}^{\pm}$ by the equalities $\partial V^{\pm}_{n}(a_{n})=a^{\pm}_{n}$, $\partial V^{\pm}_{n}(b_{n})=b^{\pm}_{n}$, $\partial V^{\pm}_{n}(c_{n})=c^{\pm}_{n}$, $\partial V^{\pm}_{n}(d_{n})=d^{\pm}_{n}$. Note that $a^{-}_{n},b^{-}_{n},c^{-}_{n},d^{-}_{n}$ are the images of $a^{+}_{n},b^{+}_{n},c^{+}_{n},d^{+}_{n}$ under $\Phi_{\Omega_{n}}$.
@@
We will show that the cross-ratio of $a^{+}_{n},b^{+}_{n},c^{+}_{n},d^{+}_{n}$ tends to 00, while the cross-ratio of $a^{-}_{n},b^{-}_{n},c^{-}_{n},d^{-}_{n}$ tends to infinity.
@@
Let $P^{+}$ be the future of $P$ as in Lemma 4.11. Let $Q$ be the timelike plane defined by the equation $x_{2}=0$. Then the path $Q\cap S^{+}_{\infty}\cap P^{+}$ from $Q\cap\overline{ab}$ to $Q\cap\overline{cd}$ along the piecewise lightlike geodesic $Q\cap S^{+}_{\infty}$ has length zero in the $\mathbb{ADS}^{2,1}$ metric. This implies that the lengths of the paths $Q\cap S_{n}^{+}\cap P^{+}$ converge to zero.
@@
Note that $S^{+}_{n}\cap P^{+}$ has a locally convex boundary with respect to the induced metric of $S_{n}^{+}$. Then the set $U^{+}_{n}:=(V^{+}_{n})^{-1}(S^{+}_{n}\cap P^{+})$ is a region of $(\mathbb{D},h_{n}^{+})$ that has a locally convex boundary, so it is globally convex. Since $a_{n},b_{n},c_{n},d_{n}$ are the intersection points of $C_{n}$ and $P^{+}$ by Proposition 4.7, the set $U^{+}_{n}$ contains the points $\partial V_{n}^{+}(a_{n})=a_{n}^{+}$, $\partial V_{n}^{+}(b_{n})=b_{n}^{+}$, $\partial V_{n}^{+}(c_{n})=c_{n}^{+}$, $\partial V_{n}^{+}(d_{n})=d_{n}^{+}$ in its ideal boundary. Also, since $U_{n}^{+}$ is convex, it contains the geodesics (with respect to the metric $h_{n}^{+}$) $\overline{a_{n}^{+}b_{n}^{+}}$ and $\overline{c_{n}^{+}d_{n}^{+}}$, which means that $S_{n}^{+}\cap P^{+}$ contains the geodesics $\gamma(\overline{a_{n}^{+}b_{n}^{+}})=V^{+}_{n}(\overline{a_{n}^{+}b_{n}^{+}})$ and $\gamma(\overline{c_{n}^{+}d_{n}^{+}})=V^{+}_{n}(\overline{c_{n}^{+}d_{n}^{+}})$.
Also note that the path $P\cap S_{n}^{+}\cap P_{+}$ crosses the two geodesics $\gamma(\overline{a_{n}^{+}b_{n}^{+}})$ and $\gamma(\overline{c_{n}^{+}d_{n}^{+}})$. Recall that the length of $P\cap S_{n}^{+}\cap P_{+}$ converges to 00, so the distance between the two geodesics $\gamma(\overline{a_{n}^{+}b_{n}^{+}})$ and $\gamma(\overline{c_{n}^{+}d_{n}^{+}})$ goes to zero. It follows that the cross-ratio of $(a^{+}_{n},b^{+}_{n},c^{+}_{n},d^{+}_{n})$ goes to 00.
@@
We apply a similar argument on the surfaces $S_{n}^{-}$. We denote by $T$ the timelike plane defined by the equation $x_{1}=0$. The path $T\cap S_{\infty}^{-}\cap P^{-}$ has length zero. By similar arguments as above (applying Lemma 4.11 on $P^{-}\cap S_{\infty}^{-}$ and considering the path $T\cap S^{-}_{\infty}\cap P^{-}$), the cross-ratio $(d_{n}^{-},a_{n}^{-},b_{n}^{-},c_{n}^{-})$ converges to 00. This implies that the cross-ratio of $(a_{n}^{-},b_{n}^{-},c_{n}^{-},d_{n}^{-})$ converges to $\infty$. We deduce that the quasi-symmetric constants of the gluing maps $\Phi_{\Omega_{n}}$ diverge.
4.3. Continuity of the gluing maps
The proof that we give for the next proposition is similar to the prof of [BDMS21, Proposition 8.2]
@@ 4.3. Continuity of the gluing maps
Up to normalizing by isometries of $\mathbb{ADS}^{2,1}$, we can assume that $\partial V^{+}_{n}(p)=(p,p)$ for any $p=0,1,\infty$. Note that after this normalization $\Omega_{n}$ converge in the Hausdorff topology to some convex subset $\Omega$ that has a quasi-circle as ideal boundary. Since we assumed that the gluing maps $\Phi_{\Omega_{n}}$ are normalized, we will also have that $\partial V^{-}_{n}(p)=(p,p)$ for any $p=0,1,\infty$.
We will show that each of the isometries $V_{n}^{\pm}:(\mathbb{D},h_{n}^{\pm})\to\partial^{\pm}\Omega_{n}$ converges to an isometry $V^{\pm}:(\mathbb{D},h^{\pm})\to\partial^{\pm}\Omega$. To do that, we will show that there exists $x_{0}\in\mathbb{D}$ such that $V^{\pm}_{n}(x_{0})$ is in a compact subset of $\mathbb{ADS}^{2,1}$, and their convergence will follow from [Sch96, Theorem 5.4].
We denote by $\Pi^{\pm}_{n,l}$ (resp $\Pi^{\pm}_{n,r}$) the left (resp the right) projection from ±$\partial^{\pm}\Omega_{n}$. Let $x_{0}\in\mathbb{D}$ be a fixed point. From Lemma 4.4, there is $A$ that depends only of $\epsilon$, $k$ and the sequence $(M_{p})_{p\in\mathbb{N}}$ such that each of $\Pi^{\pm}_{n,l}\circ V^{\pm}_{n}$ and $\Pi^{\pm}_{n,r}\circ V^{\pm}_{n}$ is an $A$ quasi-isometry from $\mathbb{H}^{2}$ to $\mathbb{H}^{2}$. Also, since $\partial V^{\pm}_{n}(p)=(p,p)$, for any $p=0,1,\infty$, the $A$ quasi-isometries $\Pi^{\pm}_{n,l}\circ V^{\pm}_{n}$ and $\Pi^{\pm}_{n,r}\circ V^{\pm}_{n}$ are normalized. Then by Lemma 4.8 there are compact subsets $K_{l}$ and $K_{r}$ of $\mathbb{H}^{2}$ such that the images $\Pi^{\pm}_{n,l}\circ V^{\pm}_{n}(x_{0})$ and $\Pi^{\pm}_{n,r}\circ V^{\pm}_{n}(x_{0})$ belong to $K_{l}$ and $K_{r}$ respectively.
Hence $V^{\pm}_{n}(x_{0})$ lie on a subset of light like geodesics $L_{x,y}$ where $x$ vary in $K_{l}$, a compact subset of $\mathbb{H}^{2}$, and $y$ vary in $K_{r}$, also a compact subset of $\mathbb{H}^{2}$. This implies that each of the sequences $V^{\pm}_{n}(x_{0})$ lies on a compact subset of $\mathbb{ADS}^{2,1}$ (recall that $\Omega_{n}$ converge in the Hausdorff sense). Then the isometries $V^{\pm}_{n}:(\mathbb{D},h^{\pm}_{n})\to\partial^{\pm}\Omega_{n}$ converge to the isometry $V^{\pm}:(\mathbb{D},h^{\pm})\to\partial^{\pm}\Omega$. Note that the gluing map $\Phi_{\Omega_{n}}$ satisfies
$\Phi_{\Omega_{n}}=\partial((\Pi^{-}_{n,l}\circ V^{-}_{n})^{-1}\circ(\Pi^{+}_{n,r}\circ V^{+}_{n}))$
and note that all maps $(\Pi^{-}_{n,l}\circ V^{-}_{n})^{-1}\circ(\Pi^{+}_{n,r}\circ V^{+}_{n})$ are normalized uniformly quasi-isometries. Since $(\Pi^{-}_{n,l}\circ V^{-}_{n})^{-1}\circ(\Pi^{+}_{n,r}\circ V^{+}_{n})$ converges to $(\Pi^{-}_{l}\circ V^{-})^{-1}\circ(\Pi^{+}_{r}\circ V^{+})$, then also $\Phi_{\Omega_{n}}$ converge to $\partial((\Pi^{-}_{l}\circ V^{-})^{-1}\circ(\Pi^{+}_{r}\circ V^{+}))=(\partial V^{-})^{-1}\circ\partial V^{+}=\Phi_{\Omega}$. Since we have also assumed that $\Phi_{\Omega_{n}}$ converges to $f$, we deduce $\Phi_{\Omega}=f$.

### v2

@@ 5.1. Definition of gluing maps
From now until the end of the paper, we will always assume that if $\Omega$ is a globally hyperbolic convex subset, then the induced metrics on its boundary components $\partial^{\pm}\Omega$ are isometric to $(\mathbb{D},h^{\pm})$, where $h^{\pm}$ have bounded derivatives of any order on the disc $\mathbb{D}$.
Note that from the proof of Lemma 5.6, the maps $\Pi_{l}^{\pm}\circ V^{\pm}$ extend to a homeomorphism :
$\partial(\Pi_{l}^{\pm}\circ V^{\pm}):\partial\mathbb{D}\to\partial_{\infty}\mathbb{H}^{2}.$
We define $\partial V^{\pm}:=(\pi^{\pm}_{l})^{-1}\circ\partial(\Pi_{l}^{\pm}\circ V^{\pm})$, where $\pi^{\pm}_{l}$ is the extension of $\Pi^{\pm}_{l}$ to the boundary at infinity.
@@
Remark 5.8.
Note that the isometries $V^{\pm}$ and the gluing map $\Phi_{\Omega}$ are defined up to composition by Möbius maps. Also note that Möbius maps do not necessarily preserve the metrics $h^{\pm}$, so the metrics $h^{\pm}$ are defined up to Möbius maps. However, we still have a uniquely defined normalized gluing map.
@@
Indeed, if $V^{+}:(\mathbb{D},h^{+})\to\partial^{+}\Omega$ and $V^{\prime+}:(\mathbb{D},h^{\prime+})\to\partial^{+}\Omega$ are two such isometries, then $(V^{\prime+})^{-1}\circ V^{+}:(\mathbb{D},h^{+})\to(\mathbb{D},h^{\prime+})$ is an isometry. In particular, since $h^{\prime+}$ and $h^{+}$ are conformal to $\mathbb{D}$, it follows that $(V^{\prime+})^{-1}\circ V^{+}$ is also conformal, that is, it belongs to $\rm{PSL}(2,\mathbb{R})$. Moreover, since by hypothesis $\partial V^{+}(p)=(p,p)$ and $\partial V^{\prime+}(p)=(p,p)$ for any $p=0,1,\infty$, it follows that $(V^{\prime+})^{-1}\circ V^{+}$ is the identity because it fixes three points on the boundary of $\mathbb{D}$.
@@
The discussion above leads to the following proposition.
@@
5.2. Properness of the map $\Phi_{\cdot}$
In this section we prove the properness of the map $\Phi_{\cdot}$ in the sense of the following proposition.
@@ 5.2. Properness of the map $\Phi_{\cdot}$
Let $\epsilon>0$ and $(M_{p})_{p\in\mathbb{N}}$ a sequence of positive numbers. For any $k>1$ there exists $k^{\prime}>1$ such that for any globally hyperbolic convex subset $\Omega$ in which the induced metrics on $\partial^{\pm}\Omega$ have curvatures in $(-\frac{1}{\epsilon},-1-\epsilon)$ and any derivative of the metric of order $p$ is bounded by $M_{p}$. We assume that the gluing map is normalized. If the gluing map $\Phi_{\Omega}$ is $k$-quasi-symmetric then $\partial_{\infty}\Omega$ is a $k^{\prime}$ quasi-circle.
@@
[BDMS21, Lemma 8.4] Let $P$ be a totally geodesic space-like plane. We denote by $P^{+}$ (resp $P^{-}$) the union of all future-oriented (resp past-oriented ) time-like geodesic segments of length $\frac{\pi}{2}$ starting orthogonally from $P$.
  • let $S\subset\mathbb{ADS}^{2,1}$ be a space-like past convex surface, and let $P$ be a space-like totally geodesic plane. In the neighborhood of all points $x\in S\cap P$ , the intersection $S\cap P^{+}$ is locally convex in the induced metric on $S$.
  • Let $S\subset\mathbb{ADS}^{2,1}$ be a space-like future convex surface, and let $P$ be a space-like totally geodesic plane. In the neighborhood of all points $x\in S\cap P$ , the intersection $S\cap P^{-}$ is locally convex in the induced metric on $S$.
@@
[…]
@@
Let’s work in the projective model of $\mathbb{ADS}^{2,1}$ with the coordinates so that in the affine chart $x_{4}=1$, $\partial_{\infty}\mathbb{ADS}^{2,1}$ is the hyperboloid $x_{1}^{2}+x_{2}^{2}=x_{3}^{2}+1$, with $\mathbb{ADS}^{2,1}$ seen as the region $x_{1}^{2}+x_{2}^{2}<x_{3}^{2}+1$. We may then arrange that the vertices of $C_{\diamond}$ are the points $(\pm\sqrt{2},0,-1)$ and $(0,\pm\sqrt{2},1)$. Since $\partial^{+}\Omega_{n}$ is in the future of $\rm{CH}(C_{n})$ but contained in $D(C_{n})$, the domain of dependence of $C_{n}$, and since both $\overline{D(C_{n})}$ (the closure of $D(C_{n})$ in $\mathbb{ADS}^{2,1}\cup\partial_{\infty}\mathbb{ADS}^{2,1}$) and $\rm{CH}(C_{n})$ converge to $\rm{CH}(C_{\diamond})$, we have that $\partial^{+}\Omega_{n}$ converge to $S^{+}_{\infty}$, where $S^{+}_{\infty}$ is the union of the two future faces of $\rm{CH}(C_{\diamond})$. Similarly, $\partial^{-}\Omega_{n}$ converge to $S^{-}_{\infty}$, where $S^{-}_{\infty}$ is the union of the two past faces of $\rm{CH}(C_{\diamond})$.
@@
Let $P$ be the space-like totally geodesic plane equal to the intersection with $\mathbb{ADS}^{2,1}$ of the plane $x_{3}=0$ in $\mathbb{R}^{3}$. Let $a=\left(-\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},0\right)$, $b=\left(-\frac{\sqrt{2}}{2},-\frac{\sqrt{2}}{2},0\right)$, $c=\left(\frac{\sqrt{2}}{2},-\frac{\sqrt{2}}{2},0\right)$, and $d=\left(\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},0\right)$. Thus, $a,b,c,d$ are the intersection points of $\partial P$ with $C_{\diamond}$ occurring in the cyclic order $a,b,c,d$.
@@
For all $n\in\mathbb{N}$, choose intersection points $a_{n},b_{n},c_{n},d_{n}$ of $C_{n}$ with $P$ such that $a_{n}\to a$, $b_{n}\to b$, $c_{n}\to c$, and $d_{n}\to d$. Such points exist since $C_{n}$ converge to $C_{\diamond}$. Define $a_{n}^{\pm},b_{n}^{\pm},c_{n}^{\pm},d_{n}^{\pm}$ by the equalities $\partial V^{\pm}_{n}(a^{\pm}_{n})=a_{n}$, $\partial V^{\pm}_{n}(b^{\pm}_{n})=b_{n}$, $\partial V^{\pm}_{n}(c^{\pm}_{n})=c_{n}$, $\partial V^{\pm}_{n}(d^{\pm}_{n})=d_{n}$. Note that $a^{-}_{n},b^{-}_{n},c^{-}_{n},d^{-}_{n}$ are the images of $a^{+}_{n},b^{+}_{n},c^{+}_{n},d^{+}_{n}$ under $\Phi_{\Omega_{n}}$.
@@
We will show that the cross-ratio of $a^{+}_{n},b^{+}_{n},c^{+}_{n},d^{+}_{n}$ tends to 0, while the cross-ratio of $a^{-}_{n},b^{-}_{n},c^{-}_{n},d^{-}_{n}$ tends to infinity.
@@
Let $P^{+}$ be the future of $P$ as in Lemma 5.14. Let $Q$ be the time-like plane defined by the equation $x_{2}=0$. Then the path $Q\cap S^{+}_{\infty}\cap P^{+}$ from $Q\cap\overline{ab}$ to $Q\cap\overline{cd}$ along the piecewise light-like geodesic $Q\cap S^{+}_{\infty}$ has length zero in the $\mathbb{ADS}^{2,1}$ metric (here, $\overline{ab}$ and $\overline{cd}$ are the geodesics inside $S_{\infty}^{+}$). This implies that the lengths of the paths $Q\cap S_{n}^{+}\cap P^{+}$ converge to zero.
@@
Note that $S^{+}_{n}\cap P^{+}$ has a locally convex boundary with respect to the induced metric of $S_{n}^{+}$. Then the set $U^{+}_{n}:=(V^{+}_{n})^{-1}(S^{+}_{n}\cap P^{+})$ is a region of $(\mathbb{D},h_{n}^{+})$ that has a locally convex boundary, so it is globally convex (because the metrics are non positively curved). Since $a_{n},b_{n},c_{n},d_{n}$ are the intersection points of $C_{n}$ and $P^{+}$ by Proposition 5.10, the set $U^{+}_{n}$ contains the points $(\partial V_{n}^{+})^{-1}(a_{n})=a_{n}^{+}$, $(\partial V_{n}^{+})^{-1}(b_{n})=b_{n}^{+}$, $(\partial V_{n}^{+}(c_{n}))^{-1}=c_{n}^{+}$, $(\partial V_{n}^{+}(d_{n}))^{-1}=d_{n}^{+}$ in its ideal boundary. Also, since $U_{n}^{+}$ is convex, it contains the geodesics (with respect to the metric $h_{n}^{+}$) $\overline{a_{n}^{+}b_{n}^{+}}$ and $\overline{c_{n}^{+}d_{n}^{+}}$, which implies that $S_{n}^{+}\cap P^{+}$ contains the geodesics $\gamma(\overline{a_{n}^{+}b_{n}^{+}}):=V_{n}^{+}(\overline{a_{n}^{+}b_{n}^{+}})$ and $\gamma(\overline{c_{n}^{+}d_{n}^{+}}):=V_{n}^{+}(\overline{c_{n}^{+}d_{n}^{+}})$ (in other words, $\gamma(\overline{a_{n}^{+}b_{n}^{+}})$ and $\gamma(\overline{c_{n}^{+}d_{n}^{+}})$ are the geodesics in $S_{n}^{+}$ that are the images of $\overline{a_{n}^{+}b_{n}^{+}}$ and $\overline{c_{n}^{+}d_{n}^{+}}$, respectively, under $V_{n}^{+}$).
[…]
@@
We apply a similar argument on the surfaces $S_{n}^{-}$. We denote by $T$ the time-like plane defined by the equation $x_{1}=0$. The path $T\cap S_{\infty}^{-}\cap P^{-}$ has length zero. By similar arguments as above (applying Lemma 5.14 on $P^{-}\cap S_{\infty}^{-}$ and considering the path $T\cap S^{-}_{\infty}\cap P^{-}$), the cross-ratio $(d_{n}^{-},a_{n}^{-},b_{n}^{-},c_{n}^{-})$ converges to 0. This implies that the cross-ratio of $(a_{n}^{-},b_{n}^{-},c_{n}^{-},d_{n}^{-})$ converges to $\infty$. We deduce that the quasi-symmetric constants of the gluing maps $\Phi_{\Omega_{n}}$ diverge.
5.3. Continuity of the map $\Phi_{\cdot}$
The proof that we give for the next proposition is similar to the proof of [BDMS21, Proposition 8.2].
@@ 5.3. Continuity of the map $\Phi_{\cdot}$
Up to normalizing by isometries of $\mathbb{ADS}^{2,1}$, we can assume that $\partial V^{+}_{n}(p)=(p,p)$ for any $p=0,1,\infty$. Note that after this normalization, $\Omega_{n}$ converges in the Hausdorff topology to some convex subset $\Omega$ that has a quasi-circle as ideal boundary. Since we assumed that the gluing maps $\Phi_{\Omega_{n}}$ are normalized, we also have that $\partial V^{-}_{n}(p)=(p,p)$ for any $p=0,1,\infty$.
We will show that each of the isometries $V_{n}^{\pm}:(\mathbb{D},h_{n}^{\pm})\to\partial^{\pm}\Omega_{n}$ converges to an isometry $V^{\pm}:(\mathbb{D},h^{\pm})\to\partial^{\pm}\Omega$. To do that, we will show that there exists $x_{0}\in\mathbb{D}$ such that $V^{\pm}_{n}(x_{0})$ lie in a compact subset of $\mathbb{ADS}^{2,1}$. Also recall that, by Proposition 5.5, the principal curvatures of $\partial^{\pm}\Omega_{n}$ are uniformly bounded. So the convergence of the isometries $V_{n}^{\pm}$ will follow from Theorem 5.3.
We denote by $\Pi^{\pm}_{n,l}$ (resp $\Pi^{\pm}_{n,r}$) the left (resp the right) projection from $\partial^{\pm}\Omega_{n}$. Let $x_{0}\in\mathbb{D}$ be a fixed point. From Lemma 5.6, there is $A$ that depends only of $\epsilon$, $k$ and the sequence $(M_{p})_{p\in\mathbb{N}}$ such that each of $\Pi^{\pm}_{n,l}\circ V^{\pm}_{n}$ and $\Pi^{\pm}_{n,r}\circ V^{\pm}_{n}$ is an $A$ quasi-isometry from $\mathbb{H}^{2}$ to $\mathbb{H}^{2}$. Also, since $\partial V^{\pm}_{n}(p)=(p,p)$, for any $p=0,1,\infty$, the $A$ quasi-isometries $\Pi^{\pm}_{n,l}\circ V^{\pm}_{n}$ and $\Pi^{\pm}_{n,r}\circ V^{\pm}_{n}$ are normalized. Then by Lemma 5.11 there are compact subsets $K_{l}$ and $K_{r}$ of $\mathbb{H}^{2}$ such that the images $\Pi^{\pm}_{n,l}\circ V^{\pm}_{n}(x_{0})$ and $\Pi^{\pm}_{n,r}\circ V^{\pm}_{n}(x_{0})$ belong to $K_{l}$ and $K_{r}$ respectively.
Hence $V^{\pm}_{n}(x_{0})$ lie on a subset of time-like geodesics $L_{x,y}$ where $x$ vary in $K_{l}$, a compact subset of $\mathbb{H}^{2}$, and $y$ vary in $K_{r}$, also a compact subset of $\mathbb{H}^{2}$. This implies that each of the sequences $V^{\pm}_{n}(x_{0})$ lies on a compact subset of $\mathbb{ADS}^{2,1}$ (recall that $\Omega_{n}$ converge in the Hausdorff sense). Then the isometries $V^{\pm}_{n}:(\mathbb{D},h^{\pm}_{n})\to\partial^{\pm}\Omega_{n}$ converge to the isometry $V^{\pm}:(\mathbb{D},h^{\pm})\to\partial^{\pm}\Omega$. Note that the gluing map $\Phi_{\Omega_{n}}$ satisfies:
$),\Phi_{\Omega_{n}}=\partial((\Pi^{-}_{n,l}\circ V^{-}_{n})^{-1}\circ(\Pi^{+}_{n,r}\circ V^{+}_{n}$)),
[…]